C02

1. Determine whether or not the set W = {(a + 2, a, 0) : a is a real number } is a subspace of R^3 with the standars operations.

2. Let V= R^2, and define addition and scalar multiplication as follows:
u+v= (u1, u2) + (v1, v2) = (u1+v1+1, u2+v2+2)
ku= k(u1,u2)=(ku1,-ku2)
Let u=(2,1) and v=(4,-1). Compute the value of 2u+v under the given operations.
(Note that V is not a vector space.)

3. Use a determinant to prove that the vectors (6,3,2), (3, 6, 0) and (0, 0, 2) form a basis of R^3.

4. Find the basis of the row space of the matrix.

1 -2 3 9
A= 0 1 3 5
0 -1 4 9
1
Expert's answer
2020-01-22T09:03:49-0500

1

"\forall x=(a+2,a,0), y=(b+2,b,0) , z=(c+2,c,0)"

Check the axioms subspace

"1) x+y=y+x\\
x+y=(a+2,a,0)+(b+2,b,0)=\\
=(a+b+4,a+b,0)\\
y+x=(b+2,b,0)+(a+2,a,0)=\\
=(b+a+4,b+a,0)\\"

Performed

"2) (x+y)+z=x+(y+z)\\
(x+y)+z=(a+b+4,a+b,0)+(c+2,c,0)=\\
=(a+b+c+6,a+b+c,0)\\
x+(y+z)=(a+2,a,0)+(b+c+4,b+c,0)=\\
=(a+b+c+6,a+b+c,0)\\"

Performed

3)

"\exist z=(t+2,t,0): x+z=x\\
x+z=(a+2,a,0)+(t+2,t,0)=\\
=(a+t+4,a+t,0)=(a+2,a,0)"

"\left \{\begin{matrix}
 a+t+4=a+2 \\
 a+t=a\\
0=0
\end{matrix}\right.\\
\left \{\begin{matrix}
 t=-2 \\
 t=0
\end{matrix}\right.\\"

The system has no solution.

W is not a subspace of "R^3"


2

"\forall u=(u_1,u_2), v=(v_1,v_2)
\\
u+v=(u_1+v_1+1,u_2+v_2+2)\\
ku=(ku_1,-ku_2)"

Then

"u=(2,1),\\ 2u=(2\cdot2,-2\cdot1)=(4,-2)\\
v=(4,-1)\\
2u+v=(4+4+1,-2+(-1)+2)=(9,-1)"


3

Find the determinant

"\begin{vmatrix}
 6 & 3&2 \\
 3& 6&0\\
0&0&2
\end{vmatrix}=\\
=6\cdot6\cdot2+3\cdot0\cdot0+3\cdot0\cdot2-\\
-2\cdot6\cdot0-3\cdot3\cdot2-0\cdot0\cdot6=72-18=54\not=0"

The vectors "(6,3,2), (3,6,0),(0,0,2)" is a basis of "R^3".


4

Find the rank of the matrix A.

"rk(A)=rk\begin{pmatrix}
 1 & -2&3&9 \\
 0 & 1&3&5\\
0&-1&4&9
\end{pmatrix}=\\
=rk\begin{pmatrix}
 1 & -2&3&9 \\
 0 & 1&3&5\\
0&0&7&14
\end{pmatrix}=3"

The basis of the row space consists of 3 vectors, for example,

"a_1=(1,-2,3,9)\\
a_2=(0,1,3,5)\\
a_3=(0,-1,4,9)"